OpenStudy (anonymous):

If guppies cost .20 less than swordtails, and three guppies and four swordtails cost $2.83, how much do guppies cost?

OpenStudy (y2o2):

i think you mean $283 , anyways , let guppies be x and swordtail be y so y -x = 20 and 3x+4y = 283 so x = $29 and y= $49

OpenStudy (anonymous):

When you solve the equation... 7xy = 283, so 283/7=40...Where did you get 29 & 49?